On a chilly Saturday morning, Roy Parker bought 30 doughnuts for $72, prepared 10 litres of hot coffee and went to a farm auction. There, he sold all the coffee and all but six dozen of the doughnuts before heading home late in the afternoon. Knowing that he could not eat six dozen doughnuts himself, and that they would be worthless by the next day, he began to consider how he might sell the remaining doughnuts. Only one possibility occurred to him. If he drove across the town and left the doughnuts in the lunchroom at the plant where he worked, he was sure that workers on the late shift would be happy to buy the doughnuts. He felt that the eight kilometre round trip to the plant would cost him about $0.25 per kilometre. It would take him about an hour to make the trip, and he decided that it would be worthwhile to drive to the plant if he were compensated $10 for his time. Roy believes that workers will 'forget' to pay for one dozen doughnuts.

Required:

Prepare an analysis to determine how much Roy would have to charge per doughnut for the doughnuts he sells at the plant in order to recover his profit goal of $10 for his time.
Answer #1

Cost of each dozen of doughnuts = $72/ 30 = $2.4 per dozen

Cost of six dozen doughnuts = $2.4 *6= $14.4

Cost of travel = 8kms * $0.25= $2

Expected profit = $10

Total selling price = $10 + $14.4+ $2 = $26.4

The cost will be recovered only for 5 doughnuts, since the  workers will 'forget' to pay for one dozen doughnuts.

Each dozen = 12 doughnuts

No of doughnuts for 5 dozens = 12*5 = 60 doughnuts

Cost per doughnut = $26.4/ 60 = $0.44